# Context is Key: A Benchmark for Forecasting with Essential Textual Information

### v1

@@ Abstract
Forecasting is a critical task in decision making across various domains. While numerical data provides a foundation, it often lacks crucial context necessary for accurate predictions. Human forecasters frequently rely on additional information, such as background knowledge or constraints, which can be efficiently communicated through natural language. However, the ability of existing forecasting models to effectively integrate this textual information remains an open question. To address this, we introduce “Context is Key” (CiK), a time series forecasting benchmark that pairs numerical data with diverse types of carefully crafted textual context, requiring models to integrate both modalities. We evaluate a range of approaches, including statistical models, time series foundation models, and LLM-based forecasters, and propose a simple yet effective LLM prompting method that outperforms all other tested methods on our benchmark. Our experiments highlight the importance of incorporating contextual information, demonstrate surprising performance when using LLM-based forecasting models, and also reveal some of their critical shortcomings. By presenting this benchmark, we aim to advance multimodal forecasting, promoting models that are both accurate and accessible to decision-makers with varied technical expertise. The benchmark can be visualized at https://servicenow.github.io/context-is-key-forecasting/v0/.
@@ 1 Introduction
The estimation of future conditions is the foundation of decision making (Hyndman & Athanasopoulos, 2018) and intelligence (Wang, 2019). Articulated as time-series forecasting, this problem pervades many domains of science and commerce. Accurate forecasting relies on several crucial decisions up to the practitioner (Hyndman & Athanasopoulos, 2018), in particular on: 1. Model selection: Choosing the appropriate forecasting model for a given problem, and 2. Incorporating prior information: Determining what relevant information should be included in the model and how to effectively integrate it. This involves decisions about statistical priors, inductive biases in the model architecture, and other forms of domain knowledge integration. Traditionally, these decisions have heavily relied on expert knowledge and manual intervention. However, recent advancements in machine learning have shown particular promise in democratizing time-series forecasting by automating both model selection and the integration of prior information.
@@
In the wake of the foundation model paradigm shift (Bommasani et al., 2021), several works (e.g.,  Liang et al. (2024); Chen et al. (2023); Lim & Zohren (2021)) have addressed automatic model selection by learning flexible, adaptable models that can be applied across various problem scenarios. Unfortunately, when compared to traditional statistical methods, current approaches provide debatable performance improvements while requiring significantly more resources (Garza & Mergenthaler-Canseco, 2024). Moreover, these models typically cast inputs and outputs as purely numerical time series, which leaves no room for the context that human experts typically rely on to focus their modelling efforts.
@@
Figure 1: An example task from the proposed Context is Key (CiK) benchmark with forecasts produced by a context-aware model. Left: Using the numerical history alone leads to poor forecasts, as nothing indicates a reversion to zero. Right: Awareness of the context significantly improves the forecasts as it reveals that no power will be produced during the night (via deductive reasoning) and enables estimating the peak hour of production.
@@
An alternative class of recent approaches (Jin et al., 2024; Liu et al., 2024b; Requeima et al., 2024) adapt large language models (LLMs) for forecasting and leverage natural language as an intuitive interface to integrate side information. These methods overcome a significant limitation of traditional forecasting techniques by eliminating the need to manually encode priors or design specialized models. They further hold the promise to capture a broader range of prior knowledge and context, potentially leading to more comprehensive and accurate forecasts. Unfortunately, there are as of yet no systematic evaluations of these models’ abilities to jointly leverage historical observations and natural language for forecasting. While several benchmarks for context-aided forecasting have been recently released (Zhang et al., 2023; Liu et al., 2024a; Xu et al., 2024; Emami et al., 2024; Merrill et al., 2024), their contexts are not guaranteed to be useful for improving performance. Hence, it is still an open question as to what extent existing methods can improve their predictions by leveraging crucially-relevant information provided in textual form.
@@
To this end, we propose the Context Is Key (CiK, pronounced kick) benchmark of forecasting tasks with numerical input-output pairs and essential textual context. The benchmark is designed to assess a forecaster’s ability to utilize both the numerical data and key information contained within the accompanying text, as the accuracy of the forecasts relies heavily on effectively leveraging both; see Fig. 1 for an example where context is imperative to forecast accuracy.
@@
  • We manually curate and release 71 forecasting tasks (Sec. 3) spanning 7 domains, which cover various kinds of contextual information (Sec. 3.2), and in addition to basic natural language-processing and time-series analysis, require various capabilities (Sec. 3.3).
  • We introduce the Region of Interest CRPS metric (RCRPS) to evaluate context-aided forecasting performance (Sec. 4), which prioritizes context-sensitive windows in the prediction and accounts for constraint satisfaction.
  • We evaluate various methods on CiK (Sec. 5), including statistical models, time series foundation models using only numerical data, and LLM-based forecasters capable of incorporating context. We introduce Direct Prompt, a simple prompting method that achieves the best results on CiK. Our analysis explores key factors such as the impact of context conditioning, prompting techniques, model capabilities, and discusses failure modes of models.
@@ 2 Problem Setting
This work addresses the problem of context-aided forecasting, where the goal is to produce statistical forecasts by incorporating relevant side information (i.e., context). Let $\mathbf{X}_{H}=[X_{1},\ldots,X_{t}]$ represent a series of random variables corresponding to historical observations, where each $X_{\tau}\in\mathcal{X}\subseteq\mathbb{R}$, and let $\mathbf{X}_{F}=[X_{t+1},\ldots,X_{T}]$ represent future observations. In the classical statistical forecasting problem, the objective is to estimate the joint distribution of the future observations given the historical data:
$P(\mathbf{X}_{F}\mid\mathbf{X}_{H}).$
We further assume access to context, denoted $\mathbf{C}\in\mathcal{C}$, which is additional data of arbitrary nature ($\mathcal{C}$) that contains information relevant for predicting $\mathbf{X}_{F}$ and complementary to the history $\mathbf{X}_{H}$. The task then becomes estimating the distribution:
$P(\mathbf{X}_{F}\mid\mathbf{X}_{H},\mathbf{C}).$
Crucially, we restrict our focus to relevant context, which we define as context that does not degrade the prediction of future time steps. Formally, for $\mathbf{x}_{F}\sim\mathbf{X}_{F}\mid\mathbf{X}_{H},\mathbf{C},$ given some loss function $\mathcal{L}$ assessing a predictive distribution over $\mathbf{X}_{F}$ against a realization $\mathbf{x}_{F}$, $\mathcal{L}:P(\mathbf{X}_{F})\times\mathbf{x}_{F}\rightarrow\mathbb{R}$, we are interested in systems where, in expectation, forecasts that leverage context perform better: [^1] Using the negative log-probability as the loss function would make this statement equivalent to: the entropy of $P(\mathbf{X}_{F}\mid\mathbf{X}_{H},\mathbf{C})$ must be lower than the cross entropy of $P(\mathbf{X}_{F}\mid\mathbf{X}_{H},\mathbf{C})$ and $P(\mathbf{X}_{F}\mid\mathbf{X}_{H})$.
@@
Furthermore, although the nature of the context $\mathcal{C}$ can vary widely, we specifically concentrate on context communicated through natural language.
@@ 3 Context is Key: a Natural Language Context-Aided Forecasting Benchmark
Figure 2: Overview: The tasks in the CiK benchmark rely on real-world numerical data, from 7 domains, as well as synthetic data (left), coupled with natural language context capturing up to 5 different aspects of the dynamical process (center), and require up to 7 non-trivial capabilities to unlock accurate forecasts (right).
We present the Context is Key (CiK) benchmark, a collection of probabilistic forecasting tasks that cannot be solved without integrating natural language contextual information with numerical data. CiK consists of 71 distinct tasks spanning seven application domains (Sec. 3.1) and that can be instantiated in different ways, e.g., by changing target time series or by selecting different time windows. These tasks encompass diverse types of contextual information (e.g., past events and known causal relationships; Sec. 3.2), and are designed such that various capabilities (e.g., causal reasoning; Sec. 3.3) are required to fully leverage the context and unlock accurate forecasts (see Fig. 2 for an overview). An example task is illustrated in Fig. 1 and others are given in Appendix B. The complete set of tasks can be explored at https://servicenow.github.io/context-is-key-forecasting/v0/ and the source code is available at https://github.com/ServiceNow/context-is-key-forecasting.
@@ 3.1 Domains and Numerical Data Sources
The vast majority (95%) of tasks in CiK draw numerical data from 2,644 real-world time series acquired from public sources. These series cover a range of domains: Climatology (solar irradiance and cloud coverage (Sengupta et al., 2018)); Economics (unemployment rates across states and counties (U.S. Bureau of Labor Statistics, 2024)); Energy (electricity consumption and production (Godahewa et al., 2021)); Mechanics (experimental properties of physical systems (Gamella et al., 2024)); Public Safety (fire department intervention counts (Ville de Montréal, 2020)); Transportation (highway segment occupancy rates and average speeds (Chen et al., 2001)); and Retail (cash withdrawals from various ATMs (Godahewa et al., 2021)). The remaining 5% of tasks use simulated data from dynamical systems crafted specifically for the tasks. Overall, the time series in CiK exhibit diverse sampling frequencies, with observations ranging from every 10 minutes to monthly intervals. Additional details on data sources can be found in Sec. A.1.
Memorization mitigation:  Using publicly available real-world data introduces the risk that pretrained LLMs and time-series foundation models may have memorized portions of the data, potentially inflating evaluation performance. To mitigate this, we employ several strategies. First, we prioritize live data sources that are continuously updated, such as Chen et al. (2001) and Ville de Montréal (2020), ensuring the data is collected after the training cut-off dates of the models being evaluated. Second, where applicable, we use derived series that are not directly available in the raw data, such as converting an incident log into time series (Ville de Montréal, 2020). Finally, as a last resort, we apply minor transformations, such as adding noise or shifting timestamps, but use these sparingly to avoid misalignment between common-sense knowledge (e.g., holiday dates) and the numerical data. Details on the mitigation methods used for each data source are provided in Sec. A.1.
@@ 3.2 Natural Language Context
[…]
For each task in the benchmark, we jointly sample numerical data from one of the series described in Sec. 3.1 and then carefully craft the natural language context necessary to unlock accurate forecasts. In some cases, this context is purely descriptive, providing information about the general nature of the target variable and its historical behavior, as seen in the task illustrated in Fig. 1. In other cases, the raw numerical data is adjusted to reflect the influence of the context. For example, in one task based on data from Godahewa et al. (2021), an ATM is expected to be inaccessible during a specific period, leading to zero withdrawals (visualized in Sec. B.3). In another task, electricity demand is projected to surge due to an incoming weather event (visualized in Sec. B.2). In such cases, we modify the series to incorporate these patterns and include such information into the corresponding context.
Overall, we include diverse forms of natural language context, capturing various aspects of the process underlying the time series and revealing complementary knowledge that could be provided by a human expert or an external information source. The types of context are described below and exemplified in the task illustrated in Fig. 3. Several additional examples are given in Appendix B.
@@
Finally, we note that, in contrast with the work of Zhang et al. (2023); Merrill et al. (2024); Liu et al. (2024a); Emami et al. (2024) which rely on LLMs or scraped news articles, all contextual information and data transformations in the CiK benchmark are manually crafted to ensure quality and relevance.
3.3 Model Capabilities
In addition to forecasting and natural language understanding, all tasks are designed such that fully utilizing the contextual information requires a range of capabilities, including instruction following, various forms of reasoning, and retrieval.
@@ 3.3 Model Capabilities
A list of all capabilities with definitions is available in Sec. A.4 and the capabilities required to solve each task are documented at https://servicenow.github.io/context-is-key-forecasting/v0/. The distributions of tasks per capability and context type are shown in Fig. 2, while the distribution of lengths of the numerical historical data, prediction horizons and natural language context can be found in Sec. A.5. Multiple example tasks from CiK are provided in Appendix B, along with an explanation of their sources of natural language context and the capabilities required to solve them.
4 Region Of Interest Continuous Ranked Probability Score
@@
Alongside the tasks, we introduce the Region of Interest CRPS (RCRPS), a novel proper scoring rule designed specifically for context-aided probabilistic forecasting. This new scoring rule is an extension of the Continuous Ranked Probability Score (CRPS; Gneiting & Raftery (2007)), a proper scoring rule that provides a comprehensive assessment of forecast quality by evaluating the entire predictive distribution rather than focusing solely on summary statistics. Importantly, since it is based on the CRPS, the RCRPS can be calculated using only samples from the predictive distribution, and so can be used even in cases where closed-form distributions are unavailable. The RCRPS extends the CRPS via two key components: a region of interest and a measure of constraint satisfaction. This allows assessing both forecast accuracy and the integration of contextual information.
@@
Region of interest (RoI):  The score reweighs a strict subset of time steps, denoted by $\mathcal{I}\subseteq[t\!+\!1,\dots,T]$, whose values are heavily informed by the context. For example, in the ATM task described in Sec. 3.2 (visualized in Sec. B.3), this would correspond to the time intervals during which the ATM is expected to be unavailable. In other tasks, such as those in Figs. 1 and 3, where the context informs the value of all future time points, we set the RoI to an empty set, essentially weighting all time steps equally (for readability, we report the definition of RCRPS for this special case in Appendix E).
@@
Constraint satisfaction:  The score penalizes violations of constraints, whether explicitly or implicitly included in the context, by measuring a task-specific function, denoted by $v_{\mathbf{C}}$, whose value is zero for any trajectory that satisfies the constraints and $>0$ for any trajectory that violates them. Concrete examples are given in Sec. E.4. For tasks whose context does not imply constraints, we use $v_{\mathbf{C}}(\cdot)\equiv 0$.
@@
Given an inferred forecast distribution $\widetilde{\mathbf{X}}_{F}$ and a ground truth $\mathbf{x}_{F}$, the scoring rule is defined as: $\text{RCRPS}(\widetilde{\mathbf{X}}_{F},\mathbf{x}_{F})\mathrel{\mathop{:}}=\mathcal{\alpha}\,\cdot\left[\frac{1}{2|\mathcal{I}|}\cdot\sum_{i\in\mathcal{I}}\text{CRPS}\!\!\left(\widetilde{X}_{i},x_{i}\right)+\frac{1}{2|\neg\mathcal{I}|}\cdot\sum_{i\in\neg\mathcal{I}}\text{CRPS}\!\!\left(\widetilde{X}_{i},x_{i}\right)+\beta\cdot\,\text{CRPS}\!\!\left(v_{\mathbf{C}}(\widetilde{\mathbf{X}}_{F}),0\right)\right],$

### v2

@@ Abstract
Forecasting is a critical task in decision-making across numerous domains. While historical numerical data provide a start, they fail to convey the complete context for reliable and accurate predictions. Human forecasters frequently rely on additional information, such as background knowledge and constraints, which can efficiently be communicated through natural language. However, in spite of recent progress with LLM-based forecasters, their ability to effectively integrate this textual information remains an open question. To address this, we introduce “Context is Key” (CiK), a time-series forecasting benchmark that pairs numerical data with diverse types of carefully crafted textual context, requiring models to integrate both modalities; crucially, every task in CiK requires understanding textual context to be solved successfully. We evaluate a range of approaches, including statistical models, time series foundation models, and LLM-based forecasters, and propose a simple yet effective LLM prompting method that outperforms all other tested methods on our benchmark. Our experiments highlight the importance of incorporating contextual information, demonstrate surprising performance when using LLM-based forecasting models, and also reveal some of their critical shortcomings. This benchmark aims to advance multimodal forecasting by promoting models that are both accurate and accessible to decision-makers with varied technical expertise. The benchmark can be visualized at https://anon-forecast.github.io/benchmark_report_dev/.
Machine Learning, ICML
[…]
@@ 1 Introduction
The prediction of future states of the world is a cornerstone of decision making (Hyndman & Athanasopoulos, 2018) and intelligence (Wang, 2019). Articulated as time-series forecasting, this problem pervades much of science and commerce.
@@
Accurate forecasting relies on several decisions up to the practitioner (Hyndman & Athanasopoulos, 2018): 1. Model selection: choosing an appropriate forecasting model for a given problem, and 2. Incorporating prior information: determining what relevant information to integrate into the model and how to do so effectively. This involves decisions about statistical priors, inductive biases in the model architecture, and other forms of domain knowledge integration, all of which traditionally rely on expert knowledge and manual intervention. However, recent advancements in machine learning have shown promise in automating both model selection and the incorporation of prior information, accelerating the democratization of time series forecasting.
@@
In the wake of the foundation model paradigm shift (Bommasani et al., 2021), several works such as  Liang et al. (2024); Chen et al. (2023); Lim & Zohren (2021), have addressed automatic model selection by learning flexible, adaptable models applicable across various problem scenarios. However, these approaches are much more costly than traditional statistical methods, and provide debatable improvements in performance (Garza & Mergenthaler-Canseco, 2024). Typically, these models process purely numerical time series, excluding the context that human forecasters rely on to incorporate prior information.
@@
Figure 1: An example task from the proposed Context is Key (CiK) benchmark with GPT-4o forecasts in blue and the ground truth in yellow. Left: Forecasts based on the numerical history alone are inaccurate, as nothing indicates a reversion to zero. Right: The context enables better forecasts because it reveals that the series represents photovoltaic power production. Hence, the model can deduce that no power will be produced at night. The context also enables better estimation of the peak hour of production by providing statistics from the history.
@@
An alternative class of recent approaches (Jin et al., 2024; Liu et al., 2024c; Requeima et al., 2024) adapt large language models (LLMs) for forecasting and leverage natural language (NL) as an intuitive interface to integrate side information. These methods overcome a significant limitation of traditional forecasting techniques by eliminating the need to manually encode priors or design specialized models. They further hold the promise of capturing a broader range of prior knowledge and context, potentially leading to more comprehensive and accurate forecasts.
Unfortunately, there are as of yet no systematic evaluations of these models’ abilities to jointly leverage historical observations and natural language for forecasting. While several benchmarks for context-aided forecasting have been recently released (Zhang et al., 2023; Liu et al., 2024a; Xu et al., 2024; Emami et al., 2024; Merrill et al., 2024), their contexts are not guaranteed to be useful for improving performance. As such, it remains unknown whether existing models can enhance their forecasts by utilizing crucially relevant textual context.
@@
To this end, we propose the Context is Key (CiK, pronounced kick) benchmark of forecasting tasks. CiK consists of tasks designed to assess a forecasting model’s ability to use both 1. numerical input-output pairs and 2. essential textual context. As shown in Figure 1, accurate forecasts in CiK are made possible only by effectively leveraging both numerical data and key information contained within the accompanying text.
@@
  • CiK Benchmark: A collection of 71 manually designed forecasting tasks spanning seven real-world domains, each requiring the integration of diverse contextual information that has a non-trivial impact on forecasts (Sec. 3).
  • Region of Interest CRPS (RCRPS): A scoring rule to evaluate context-aided forecasting performance, which prioritizes context-sensitive windows and accounts for constraint satisfaction (Sec. 4).
  • Direct Prompt Forecasters: A simple yet effective prompt-based approach to using LLMs as context-aided forecasters, which serves as a surprisingly strong baseline on CiK (Sec. 5.2).
  • Extensive evaluation of diverse models on CiK, including statistical models, time series foundation models using only numerical data, and LLM-based forecasters capable of incorporating natural language context. Our analysis explores key factors such as the impact of context conditioning, tradeoffs in model size and performance, and discusses failure modes of models (Sec. 5).
@@ 2 Problem Setting
This work addresses the problem of context-aided forecasting, where the goal is to produce statistical forecasts by incorporating relevant side information (i.e., context). Let $\mathbf{X}_{H}=[X_{1},\ldots,X_{t}]$ represent a series of random variables corresponding to historical observations in discrete time, where each $X_{\tau}\in\mathcal{X}\subseteq\mathbb{R}$, and let $\mathbf{X}_{F}=[X_{t+1},\ldots,X_{T}]$ represent future observations. In classical statistical forecasting, the goal is to estimate the joint distribution of future observations given the historical ones, $P(\mathbf{X}_{F}\mid\mathbf{X}_{H}).$ We further assume access to context, denoted $\mathbf{C}$, which consists of additional data of arbitrary nature containing information relevant for predicting $\mathbf{X}_{F}$ and complementary to the history $\mathbf{X}_{H}$. The task then becomes estimating the distribution $P(\mathbf{X}_{F}\mid\mathbf{X}_{H},\mathbf{C}).$ Crucially, we restrict our focus to relevant context, which we define as context that does not degrade forecasts. Formally, for $\mathbf{x}_{F}\sim\mathbf{X}_{F}\mid\mathbf{X}_{H},\mathbf{C},$ given some loss function $\mathcal{L}$ assessing a predictive distribution over $\mathbf{X}_{F}$ against a realization $\mathbf{x}_{F}$, $\mathcal{L}:P(\mathbf{X}_{F})\times\mathbf{x}_{F}\rightarrow\mathbb{R}$, we are interested in models where, in expectation, forecasts that leverage context perform better:
@@
Furthermore, although the nature of the context can vary widely, we specifically concentrate on context communicated through natural language, which we refer to as “context” or “contextual information”.
@@ 3 Context is Key: a Natural Language Context-Aided Forecasting Benchmark
Figure 2: The tasks in the CiK benchmark rely on real-world numerical data from 7 domains.
We introduce the Context is Key (CiK) benchmark, a collection of probabilistic forecasting tasks where accurate predictions require integrating both numerical data and natural language contextual information. CiK comprises 71 distinct tasks spanning seven real-world application domains (see Sec. 3.1), each featuring various stochastic components that can be instantiated into thousands of task instances (e.g., time series, time windows, natural language formulation). These tasks encompass diverse types of contextual information that reveal various aspects of dynamical processes (see Sec. 3.2). Moreover, they are designed such that context is key in that it non-trivially unlocks accurate forecasts, e.g., by conveying causal relationships that reveal the effect of a covariate on the time series of interest. An example task is shown in Figure 1 and others can be found in Appendix B.
Availability: CiK is open source. The complete set of tasks can be explored at https://bit.ly/cik_tasks and the source code, at https://bit.ly/cik_code. All data sources used for CiK’s tasks are openly available (see Sec. A.1).
@@ 3.1 Domains and Numerical Data Sources
As illustrated in Figure 2, the majority (95%) of tasks in CiK are based on real-world application domains. We leverage 2,644 time series sourced from publicly available datasets across seven domains: Climatology (solar irradiance and cloud coverage (Sengupta et al., 2018)); Economics (unemployment rates across states and counties (U.S. Bureau of Labor Statistics, 2024)); Energy (electricity consumption and production (Godahewa et al., 2021)); Mechanics (experimental properties of physical systems (Gamella et al., 2024)); Public Safety (fire department intervention counts (Ville de Montréal, 2020)); Transportation (highway segment occupancy rates and average speeds (Chen et al., 2001)); and Retail (cash withdrawals from various ATMs (Godahewa et al., 2021)). The remaining 5% of tasks use simulated data from dynamical systems crafted specifically for the tasks. Overall, the time series in CiK exhibit diverse sampling frequencies, with observations ranging from every 10 minutes to monthly intervals; additional details on data sources can be found in Sec. A.1.
Memorization mitigation:  Building tasks with publicly available data introduces contamination risk: pretrained LLMs and time-series foundation models may have memorized portions of the data, potentially inflating evaluation performance. We employ several mitigation strategies. First, we prioritize live data sources that are continuously updated, such as Chen et al. (2001) and Ville de Montréal (2020), ensuring the data is collected after the training cut-off dates of the models that we evaluate. Second, where applicable, we use derived series that are not directly available in the raw data, such as incident logs converted into time series (Ville de Montréal, 2020). Finally, as a last resort, we apply minor transformations, such as adding noise or shifting timestamps, but use these sparingly to avoid misalignment between common-sense knowledge (e.g., holiday dates) and the numerical data. The exact mitigation methods used, per data source, are given in Sec. A.1.
@@ 3.2 Natural Language Context
Figure 3: Number of tasks per context type in CiK.
For each task in the benchmark, we jointly sample numerical data from one of the series described in Sec. 3.1 and then manually craft the natural language context necessary to unlock accurate forecasts. In some cases, this context is purely descriptive, providing information about the general nature of the target variable and its historical behavior, as seen in the task illustrated in Figure 1. In other cases, the raw numerical data is adjusted to reflect the influence of the context. For example, in one task based on data from Godahewa et al. (2021), an ATM is expected to be inaccessible during a specific time period in the future, leading to zero withdrawals (see Sec. B.3). In another task, electricity demand is projected to surge due to an incoming weather event (see Sec. B.2). For such cases, we modify the series to incorporate patterns described by the context.
Overall, we include diverse forms of natural language context, each capturing a different aspect of the process underlying the time series and providing complementary knowledge that a human expert could leverage for more accurate forecasting. The types of context are described below and exemplified in the task illustrated in Figure 4. For additional clarity, further examples are provided in Appendix B and the distribution of tasks per context type is shown in Figure 3.
@@
Finally, for completeness, Sec. A.7 provides the distributions of lengths of the numerical historical data, prediction horizons and natural language context.
@@
3.3 Validating the Relevance of The Context
Related efforts on context-aided forecasting, outlined in Sec. 6, rely on scraping and/or LLMs to obtain natural language context (Zhang et al., 2023; Merrill et al., 2024; Liu et al., 2024a; Emami et al., 2024). In contrast, to ensure both the quality and relevance of tasks in CiK, we manually craft all contextual information and associated data transformations according to the procedure described in Sec. A.2. To validate the importance of the context, we subject each task to review by a panel of human and LLM evaluators tasked with assessing if the context enables better forecasts. The results are overwhelmingly positive, with humans finding the context to be relevant for $95\%$ of evaluated instances (details in Sec. A.4).
4 Region of Interest CRPS
@@
Alongside the tasks, we introduce the Region of Interest Continuous Ranked Probability Score (RCRPS), a novel proper scoring rule designed specifically for context-aided probabilistic forecasting. This new scoring rule is an extension of the Continuous Ranked Probability Score (CRPS; Gneiting & Raftery (2007)), a proper scoring rule that provides a comprehensive assessment of forecast quality by evaluating the entire predictive distribution rather than point forecasts. Since it is based on the CRPS, the RCRPS can be calculated using only samples from the predictive distribution, and so can be used even in cases where closed-form distributions are unavailable. The RCRPS extends the CRPS via two key components: a region of interest and a measure of constraint satisfaction.
@@
Region of interest (RoI):  The RCRPS assigns more weight to errors in a task’s RoI, which is a subset of time steps $\mathcal{I}\subseteq[t\!+\!1,\dots,T]$ for which the context is particularly relevant. For example, in the ATM task from Sec. 3.2 (visualized in Sec. B.3), the RoI denotes the time steps during which the ATM is expected to be unavailable. In other tasks, such as those in Figures 1 and 4, where the context informs the value of all future time points, the RCRPS assigns equal weights to all time steps (for readability, we report the definition of RCRPS for this special case in Appendix E).
Table 1: Results of selected models on the CiK benchmark. Starting from the left, the first column shows the RCRPS averaged over all tasks. The second column shows the rank of each method w.r.t. other models, averaged over all tasks. The remaining columns show the average RCRPS stratified by types of context (Sec. 3.2). All averages are weighted according to the scheme described in Sec. 5.1 and accompanied by standard errors. Lower is better and the best averages are in bold. An asterisk (*) denotes models that do not use natural language context. For results on all models and with alternative aggregation strategies, see Appendix C.
Model Average RCRPS Average Rank Intemporal Information Historical Information Future Information Covariate Information Causal Information
Direct Prompt (ours)
Llama-3.1-405B-Inst
0.159 $\pm$ 0.008 4.447 $\pm$ 0.237 0.174 $\pm$ 0.010 0.146 $\pm$ 0.001 0.066 $\pm$ 0.004 0.168 $\pm$ 0.010 0.398 $\pm$ 0.045
Llama-3-70B-Inst
0.286 $\pm$ 0.004 7.749 $\pm$ 0.108 0.336 $\pm$ 0.006 0.180 $\pm$ 0.003 0.185 $\pm$ 0.006 0.230 $\pm$ 0.005 0.629 $\pm$ 0.019
Mixtral-8x7B-Inst
0.523 $\pm$ 0.023 14.478 $\pm$ 0.149 0.724 $\pm$ 0.037 0.236 $\pm$ 0.002 0.185 $\pm$ 0.006 0.384 $\pm$ 0.028 0.875 $\pm$ 0.128
Qwen-2.5-7B-Inst
0.289 $\pm$ 0.003 11.273 $\pm$ 0.250 0.289 $\pm$ 0.004 0.176 $\pm$ 0.003 0.281 $\pm$ 0.007 0.238 $\pm$ 0.002 0.525 $\pm$ 0.003
Qwen-2.5-0.5B-Inst
0.464 $\pm$ 0.012 12.755 $\pm$ 0.173 0.610 $\pm$ 0.019 0.165 $\pm$ 0.004 0.221 $\pm$ 0.011 0.483 $\pm$ 0.016 0.429 $\pm$ 0.006
GPT-4o
0.276 $\pm$ 0.010 4.519 $\pm$ 0.156 0.220 $\pm$ 0.007 0.118 $\pm$ 0.001 0.113 $\pm$ 0.001 0.258 $\pm$ 0.012 0.858 $\pm$ 0.053
GPT-4o-mini
0.353 $\pm$ 0.022 9.032 $\pm$ 0.193 0.474 $\pm$ 0.035 0.139 $\pm$ 0.002 0.139 $\pm$ 0.002 0.343 $\pm$ 0.029 0.644 $\pm$ 0.128
LLMP
Llama-3-70B-Inst
0.550 $\pm$ 0.013 8.342 $\pm$ 0.234 0.455 $\pm$ 0.018 0.516 $\pm$ 0.028 0.751 $\pm$ 0.022 0.573 $\pm$ 0.017 0.392 $\pm$ 0.028
Llama-3-70B
0.237 $\pm$ 0.006 6.549 $\pm$ 0.245 0.213 $\pm$ 0.005 0.121 $\pm$ 0.008 0.282 $\pm$ 0.016 0.194 $\pm$ 0.004 0.360 $\pm$ 0.011
Mixtral-8x7B-Inst
0.264 $\pm$ 0.004 8.475 $\pm$ 0.261 0.242 $\pm$ 0.007 0.173 $\pm$ 0.004 0.312 $\pm$ 0.012 0.217 $\pm$ 0.002 0.437 $\pm$ 0.007
Mixtral-8x7B
0.262 $\pm$ 0.008 8.528 $\pm$ 0.198 0.250 $\pm$ 0.008 0.119 $\pm$ 0.003 0.297 $\pm$ 0.018 0.227 $\pm$ 0.007 0.457 $\pm$ 0.011
Qwen-2.5-7B-Inst
1.980 $\pm$ 0.027 18.475 $\pm$ 0.276 2.518 $\pm$ 0.044 2.857 $\pm$ 0.056 1.891 $\pm$ 0.008 1.605 $\pm$ 0.036 1.333 $\pm$ 0.144
Qwen-2.5-7B
0.904 $\pm$ 0.037 15.976 $\pm$ 0.343 1.140 $\pm$ 0.047 1.002 $\pm$ 0.053 0.594 $\pm$ 0.068 0.633 $\pm$ 0.048 0.928 $\pm$ 0.129
Qwen-2.5-0.5B-Inst
1.938 $\pm$ 0.024 20.091 $\pm$ 0.195 2.445 $\pm$ 0.038 1.960 $\pm$ 0.063 1.563 $\pm$ 0.009 1.762 $\pm$ 0.031 1.199 $\pm$ 0.129
Qwen-2.5-0.5B
1.991 $\pm$ 0.024 19.691 $\pm$ 0.272 2.539 $\pm$ 0.039 2.083 $\pm$ 0.052 1.651 $\pm$ 0.013 1.796 $\pm$ 0.031 1.225 $\pm$ 0.128
Multimodal Models
UniTime
0.370 $\pm$ 0.001 14.657 $\pm$ 0.091 0.457 $\pm$ 0.002 0.155 $\pm$ 0.000 0.202 $\pm$ 0.003 0.396 $\pm$ 0.001 0.423 $\pm$ 0.001
Time-LLM (ETTh1)
0.476 $\pm$ 0.001 17.941 $\pm$ 0.076 0.517 $\pm$ 0.002 0.183 $\pm$ 0.000 0.400 $\pm$ 0.002 0.443 $\pm$ 0.001 0.482 $\pm$ 0.001
TS Foundation Models*
Lag-Llama
0.329 $\pm$ 0.004 13.419 $\pm$ 0.235 0.333 $\pm$ 0.005 0.167 $\pm$ 0.005 0.281 $\pm$ 0.008 0.299 $\pm$ 0.004 0.495 $\pm$ 0.014
Chronos-Large
0.326 $\pm$ 0.002 12.295 $\pm$ 0.149 0.314 $\pm$ 0.002 0.179 $\pm$ 0.003 0.360 $\pm$ 0.003 0.253 $\pm$ 0.002 0.460 $\pm$ 0.004
TimeGEN
0.354 $\pm$ 0.000 15.075 $\pm$ 0.096 0.333 $\pm$ 0.000 0.177 $\pm$ 0.000 0.385 $\pm$ 0.000 0.292 $\pm$ 0.000 0.474 $\pm$ 0.000
Moirai-Large
0.520 $\pm$ 0.006 12.857 $\pm$ 0.260 0.596 $\pm$ 0.009 0.140 $\pm$ 0.001 0.414 $\pm$ 0.002 0.505 $\pm$ 0.007 0.438 $\pm$ 0.011
Statistical Models*
ARIMA
0.480 $\pm$ 0.006 12.740 $\pm$ 0.165 0.565 $\pm$ 0.010 0.200 $\pm$ 0.007 0.332 $\pm$ 0.003 0.386 $\pm$ 0.006 0.440 $\pm$ 0.011
ETS
0.522 $\pm$ 0.009 14.962 $\pm$ 0.199 0.627 $\pm$ 0.014 0.362 $\pm$ 0.014 0.295 $\pm$ 0.005 0.397 $\pm$ 0.010 0.508 $\pm$ 0.017
Exp-Smoothing
0.603 $\pm$ 0.013 15.673 $\pm$ 0.152 0.700 $\pm$ 0.020 0.493 $\pm$ 0.016 0.378 $\pm$ 0.006 0.485 $\pm$ 0.016 0.827 $\pm$ 0.060
@@
Constraint satisfaction:  The RCRPS penalizes constraint violations according to a task-specific function $v_{\mathbf{C}}$ whose value is positive for any trajectory that violates the constraints. Concrete examples are given in Sec. E.4.
@@
Given an inferred forecast distribution $\widetilde{\mathbf{X}}_{F}$ and a ground truth $\mathbf{x}_{F}$, the scoring rule is defined as:
$\text{RCRPS}(\widetilde{\mathbf{X}}_{F},\mathbf{x}_{F})\;\mathrel{\mathop{:}}=\;\alpha\,\cdot\Biggl{[}\frac{1}{2|\mathcal{I}|}\sum_{i\in\mathcal{I}}\text{CRPS}\!\bigl{(}\widetilde{X}_{i},x_{i}\bigr{)}+\\ \qquad\qquad\frac{1}{2|\neg\mathcal{I}|}\sum_{i\in\neg\mathcal{I}}\text{CRPS}\!\bigl{(}\widetilde{X}_{i},x_{i}\bigr{)}\;+\;\beta\,\cdot\,\text{CRPS}\!\bigl{(}v_{\mathbf{C}}(\widetilde{\mathbf{X}}_{F}),0\bigr{)}\Biggr{]},$